OpenStudy (anonymous):

sue is buying cans of bean she has several choices. select the best buy from the choices below: a) 6 cans for $2.28 b)a half case (12 cans) for $4.20 c.)40 cents per can d.) 3 cans for $1.00

OpenStudy (anonymous):

a) 2.28/6 = $0.38 per can b) 4.20/12 = $0.35 per can c) $0.40 per can d) 1/3 = $0.333333 = $0.33 per can D is the most cost efficient.